OpenStudy (anonymous):

I am having a difficult time with Radical Equations. Here is one of the problems: 4 square root x) +7 = 23. If someone could show me step by step how to do these types of problems I would be grateful. :3

OpenStudy (dumbcow):

\[4\sqrt{x} +7 = 23\] subtract 7 \[4\sqrt{x} = 16\] divide 4 \[\sqrt{x} = 4\] square both sides \[(\sqrt{x})^{2} = 4^{2}\] squaring cancels out a square root \[x = 16\]

OpenStudy (anonymous):

Thank you so much :D This helped a bunch <3